OpenStudy (anonymous):

find slope and y intercept of the line 5x=6y+48

OpenStudy (saifoo.khan):

5x - 6y = 48 Now u should too.

OpenStudy (aroub):

or yea try solving it!

OpenStudy (anonymous):

5x=6y+48 rearranging to 6y=5x-48 divide by 6 y=(5/6)x-8 from here you can see that slope=5/6 ans when x=0, y=-8, (y-intercept) ans